# CLIP-UP: A Simple and Efficient Mixture-of-Experts CLIP Training Recipe with Sparse Upcycling

### v1

@@ Abstract
Machine Learning, ICML
Apple
@@ 1 Introduction
CLIP (Radford et al., 2021; Jia et al., 2021) has proven to be a transformative model across a wide range of domains, including image classification, multimodal retrieval, and AI-driven multimodality content generation (Zhou et al., 2022; Rao et al., 2022; Gan et al., 2022; Ramesh et al., 2021; Liu et al., 2023). However, as demands on CLIP grow across these domains, scaling up the model becomes increasingly important to achieve better performance. Recent efforts to scale up CLIP have primarily focused on increasing the size of dense models (Cherti et al., 2023), which, while effective, result in significant computational overhead and high inference costs.
Figure 1: Our proposed Mixture-of-Experts (MoE) CLIP pre-training recipe. We identify several key ingredients for efficient and effective training including shared vs. separated image-text backbones, training from scratch vs. sparse upcycling from a dense model, use of auxiliary losses, and etc. A detailed analysis is provided in Section 4.1. We conclude a simple strategy that outperforms previous practice (Mustafa et al., 2022) with even less overall training cost.
Figure 2: CLIP-UP overview (Left) and sparse upcycling initialization (Right). To upcycle from a pre-trained dense checkpoint (e.g., single expert), the MLP layers in some transformer blocks are replaced with MoE layers. During initialization, the MoE MLP parameters are copied from the original dense checkpoint. The MoE router is randomly initialized.
Another promising approach to scaling CLIP is through sparse modeling, particularly by introducing Mixture-of-Expert (MoE) layers (Mustafa et al., 2022). MoE is a technique that activates only a subset of the model’s experts (parameters) per input, which reduces inference cost (Shazeer et al., 2017) compared to the dense model with the same number of parameters. MoE models have demonstrated great success in various tasks, including multimodal tasks (Fedus et al., 2022; Lepikhin et al., 2020). Despite these advances, there remains key problems in prohibitively high training cost of CLIP like models from scratch and the need of complicated auxiliary losses to stabilize the MoE model training. Mustafa et al. (2022) investigated training the LIMOE model – an MoE CLIP model with a shared backbone for text and image encoders – from scratch. While it significantly outperforms the dense model, LIMOE demands substantial computational resources and introduces newly defined auxiliary losses for enhanced robustness. For instance, LIMOE with CLIP B/16 requires 1.35 times the training FLOPs of its dense counterpart.
In this work, we explore an alternative training strategy: upcycling a pre-trained dense CLIP model to a sparse MoE architecture. Sparse upcycling (Komatsuzaki et al., 2023) leverages a well-trained dense model to initialize a sparse model with MoE layers. Figure 2 illustrates the process in detail. Our experiments start with the CLIP-B/16 architecture from LIMOE, initially trained from scratch with a shared backbone. We then further evaluate the impact of LIMOE auxiliary loss, shared versus separated backbones, and training from scratch versus sparse upcycling. As shown in Figure 1, sparse upcycling with a separated backbone yields the best quality metrics while significantly reducing training costs, lowering ZFLOPS from 4.2 to 3.7 compared to training from scratch. Incorporating LIMOE’s local and global entropy loss  (Mustafa et al., 2022) substantially improves the performance of shared backbone trained from scratch but still lags behind other configurations. Section 4.1 provides a detailed analysis of training strategies and the impact of LIMOE auxiliary loss under different setups.
In particular, we propose CLIP-UP, a sparse upcycling approach (Komatsuzaki et al., 2023) for CLIP models with MoE. As illustrated in Figure 2, this method upgrades a pretrained CLIP into a sparsely activated MoE model with relatively low additional training cost. By leveraging the pretrained dense model, CLIP-UP benefits from a warm start, significantly improving training efficiency. We demonstrate that CLIP-UP outperforms both continued training of the original dense model and training a sparse model from scratch across different scales.
In summary, our contributions are as follows:
@@
    We introduce CLIP-UP, a simple yet effective training recipe for Mixture-of-Experts (MoE) CLIP models leverages sparse upcycling from pre-trained weights, eliminating the need for complex auxiliary losses. CLIP-UP outperforms existing MoE CLIP training methods, regardless of whether the backbone is shared or separated.
@@
    CLIP-UP achieves substantial performance improvements across various Text-Image retrieval benchmarks compared to the corresponding dense models. Notably, CLIP-UP with a B/16 backbone surpasses dense CLIP by 7.2% and 5.5% on recall@1 of COCO (Lin et al., 2014) and Flickr30K (Plummer et al., 2017) text-to-image retrieval tasks respectively.
@@
    We show CLIP-UP is scalable to different model sizes (from B/32 to L/14) and provide a comprehensive analysis of key factors and challenges, providing valuable insights into the design decisions.
2 Related Work
Mixture-of-Experts. Sparsely-activated MoE models have emerged as a powerful approach for scaling model architectures while maintaining computational efficiency (Shazeer et al., 2017; Fedus et al., 2022). By activating only a subset of experts during inference, MoE models significantly reduce the required computation compared to dense models, without sacrificing performance (Jiang et al., 2024; Dai et al., 2024; Xue et al., 2024). However, the additional expert layers and the complexity of routing mechanisms required to select active experts during training make MoE models more resource-intensive to train from scratch.
Early pioneer work like LIMOE (Mustafa et al., 2022) applies the MoE framework to CLIP models, using a shared structure for both text and image modalities. This design enables efficient learning of shared representations, leading to significant performance improvements across various multimodal tasks and outperforming compute-matched dense models.
Despite its advantages, LIMOE shares a common limitation with other MoE models: the need for substantial computational resources during training (Du et al., 2024). As shown in Figure 1, training with a shared backbone from scratch yields suboptimal results compared to alternative backbone selections and training strategies. While LIMOE auxiliary loss offers some improvement, the overall performance remains inferior, suggesting that the methodology may not be the most effective approach.
Sparse upcycling. Sparse upcycling (Komatsuzaki et al., 2023) is a technique that effectively reduce training cost of sparse models by reusing existing dense model checkpoints. By leveraging the prior knowledge from a well-trained dense model, the sparse version benefits from accelerated convergence and reduced computational overhead (He et al., 2024).
The current applications of sparse upcycling have been primarily focused on single-modality models (Komatsuzaki et al., 2023; He et al., 2024). The effectiveness of this approach in the context of multimodal models remains largely unexplored. The concurrent work CLIP-MoE (Zhang et al., 2024) upcycles MoE layers by initializing them with fine-tuned MLP layers via cluster-and-contrast learning. Each expert requires an additional training stage. Scaling this method is challenging, as adding more experts increases complexity and necessitates additional training stages.
In this work, we extend sparse upcycling to CLIP with MoE using a single-stage training process. This approach significantly reduces training complexity and cost while achieving superior results, like shown in Figure 1. We also demonstrate its generalizability across shared or separated backbones and model scales. This establishes sparse upcycling as a practical and scalable solution for building efficient, high-performance CLIP models.
3 Sparse Upcycling with Mixture-of-Experts
Figure 2 outlines the CLIP-UP architecture and its upcycle training strategy. The left panel showcases the traditional CLIP contrastive training with expert layers replacing some MLP layers. The right panel demonstrates the innovative upcycle training approach. We’ll dive into the specifics of each in this section.
@@ 3.1 CLIP
Given $n$ pairs of image and text captions $\{(\mathbf{I}_{j},\mathbf{T}_{j})\}_{j=1}^{n}$, CLIP (Radford et al., 2021) learns image representations as $f(\mathbf{I}_{j})$ and text representation as $g(\mathbf{T}_{j})$ via contrastive learning objective within a training batch. Relevant pairs are positioned closer in the feature space, while irrelevant pairs are pushed farther apart. Given batch size of $B$ and a learned temperature parameter $\theta$, the image to text loss is
$\mathcal{L}_{I-T}=-\frac{1}{B}\sum_{j=1}^{B}\log\frac{e^{\text{sim}(f(\mathbf{I}_{j}),g(\mathbf{T}_{j}))/\theta}}{\sum_{k=1}^{B}e^{\text{sim}(f(\mathbf{I}_{j}),g(\mathbf{T}_{k}))/\theta}}$
and text to image loss is
$\mathcal{L}_{T-I}=-\frac{1}{B}\sum_{j=1}^{B}\log\frac{e^{\text{sim}(f(\mathbf{I}_{j}),g(\mathbf{T}_{j}))/\theta}}{\sum_{k=1}^{B}e^{\text{sim}(f(\mathbf{I}_{k}),g(\mathbf{T}_{j}))/\theta}}$
The contrastive loss is calculated by averaging the above two losses
$\mathcal{L}_{\text{Contrastive}}=\frac{1}{2}(L_{I-T}+L_{T-I})$
@@ 3.2 CLIP with Mixture-of-Experts upcycling
In CLIP-UP, for half of the transformer layers, we replace the dense MLPs with sparsely-gated MoE layers as an alternating dense-sparse-dense structure, similar to Figure 2 (Left). This helps the stability of feedforwarding and prevents gradient oscillation, especially during initial training.
@@
Each MoE layer contains $E$ experts and a router. The experts are MLPs, selectively activated based on the input. The router selects the top-$K$ experts from $E$ total experts, by predicting normalized gating logits for each input token. Let $\mathbf{X}_{j}\in\mathbb{R}^{D}$ denotes the input of $j$-th token to an MoE layer, $\mathbf{G}_{j}\in\mathbb{R}^{D\times E}$ is the gating logis for $E$ experts, $\mathbf{W}_{e}\in\mathbb{R}^{D}$ is the router weight matrix for expert $e$, the output $MoE(\mathbf{X}_{j})$ is computed as follows:
$MoE(\mathbf{X}_{j})=\mathbf{X}_{j}+\sum_{e\in\text{Top-K}}g_{e,j}MLP_{e}(\mathbf{X}_{j})$
$g_{e,j}=\begin{cases}\mathbf{G}_{e,j},&\mathbf{G}_{e,j}\in\text{Top-K}({\mathbf{G}_{e,j}|1\leq e\leq E},K),\\ 0,&\text{otherwise},\end{cases}$
$\mathbf{G}_{e,j}=Softmax(\mathbf{W}_{e}^{T}\mathbf{X}_{j})$
Each expert is assigned a fixed buffer capacity (Fedus et al., 2022), meaning it can process only a limited number of tokens at a time. With predefined capacity factor $C$, and $B_{t}$ denotes the number of tokens per batch, the capacity size of each expert $B_{e}$ is computed as $B_{e}=(B_{t}/E)\times C$. This constraint enhances computational efficiency by preventing experts from being overwhelmed, while also ensuring that hardware and memory resources are managed effectively.
Tokens are assigned to experts on a ”first-come-first-serve” basis, meaning they are processed in the order they arrive (Fedus et al., 2022). This simple and effective mechanism handles input streams effectively without introducing additional complexity or overhead from token prioritization or sorting, ensuring that the system remains efficient while distributing tokens across experts.
Auxiliary loss. While the simplified token assignment reduces overhead, it presents the challenge of maintaining a balanced token distribution across all experts. Poor distribution can result in significant token dropping (Zeng et al., 2024), which can severely impact performance. Therefore, achieving balance is essential to avoid overloading or underutilizing certain experts. Overuse of some experts may lead to overfitting, while others may not receive enough data for effective training.
To mitigate this, we follow (Zoph et al., 2022) to introduce an auxiliary loss comprising load balance loss and router $z$-loss for each modality, promoting even token distribution. Specifically, the load balance loss encourages the gating mechanism to allocate tokens more uniformly over $E$ total experts. For a sequence of length $S$, the load balance loss is defined as:
$\mathcal{L}_{\text{Balance}}=\sum_{e=1}^{E}R_{e}\cdot P_{e}$
$R_{e}=\frac{E}{K\cdot S}\sum_{j=1}^{S}\mathbbm{1}(\text{Token $j$ dispatched to Expert $e$})$
$P_{e}=\frac{1}{S}\sum_{j=1}^{S}\mathbf{G}_{e,j}$
@@
where $R_{e}$ represents the token assignment ratio to expert $e$, and $P_{e}$ denotes the average router probability of expert $e$.
@@
The router $z$-loss is to stabilize the gating process by regularizing the output of the router logits, and it ensures the router outputs stay in a reasonable range. It is defined as:
@@
$\mathcal{L}_{\text{Router}}=\frac{1}{S}\sum_{j=1}^{S}\left(\log\sum_{e=1}^{E}e^{\mathbf{G}_{e,j}}\right)^{2}$
The final training loss is the sum of the contrastive loss and auxiliary loss, with $\alpha$ and $\beta$ as scaling factors,
$\mathcal{L}=\mathcal{L}_{\text{Contrastive}}+\alpha\cdot\mathcal{L}_{\text{Balance}}+\beta\cdot\mathcal{L}_{\text{Router}}$
LIMOE auxiliary loss. LIMOE (Mustafa et al., 2022) proposes two new auxiliary losses, the local entropy loss and the global entropy loss, which are applied on a per-modality basis. The local entropy loss is defined as
$\mathcal{L}_{\text{Local}}=\frac{1}{S}\sum_{j=1}^{S}\left(-\sum_{e=1}^{E}\mathbf{G}_{e,j}\log\mathbf{G}_{e,j}\right)$
This encourages concentrated router weights but may reduce the diversity of expert utilization. The global entropy loss is defined as
$\mathcal{L}_{\text{Global}}=-\sum_{e=1}^{E}P_{e}\log P_{e}$
which promotes a balanced expert distribution, and enhance diversity in expert usage. To prevent a single modality from overusing too many experts, the global entropy loss is thresholded as:
$\mathcal{L}_{\text{Global}}^{\tau}=\max\{0,\tau+\mathcal{L}_{\text{Global}}\}$
where $\tau=\log(m)$, ensuring that at least $m$ experts are utilized to minimize the loss. These auxiliary losses stabilize training and improve performance, as demonstrated by (Mustafa et al., 2022).
According to the best recipe of LIMOE, applying the total training loss becomes:
$\mathcal{L}=$ $\mathcal{L}_{\text{Contrastive}}+\alpha\cdot\mathcal{L}_{\text{Balance}}+\beta\cdot\mathcal{L}_{\text{Router}}$ (11)
$+\gamma\cdot\mathcal{L}^{\text{Text}}_{\text{Local}}+\lambda\cdot\mathcal{L}^{\tau,\text{Image}}_{\text{Global}}+\delta\cdot\mathcal{L}^{\tau,\text{Text}}_{\text{Global}}$
In our experiments, we replaced the auxiliary loss with the LIMOE training recipe and tuned the hyperparameters for the best perfomrance. While the results demonstrate that the LIMOE auxiliary loss significantly enhances the performance of the shared backbone trained from scratch, its performance still lags behind other configurations and the same recipe doesn’t provide similar improvements in those settings. So we chose equation 3.2 as our final loss. More details will be illustrated in Section 4.
@@
Table 1: Ablation study of combinations of shared vs. separated backbone and training from scratch vs. sparse upcycling. We benchmark the models on ImageNet Zero-shot classification (Accuracy@1 %); COCO and Flickr30k text-to-image (T2I) / image-to-text (I2T) retrieval (Recall@1 %).
@@
3.3 Sparse Upcycling Training
The sparse upcycling process starts from a pre-trained dense CLIP model. The MoE encoders in the upcycled CLIP follow the architecture of the transformer used in (Komatsuzaki et al., 2023) including the number of layers and feature size, similarly for our separated backbone setup. A selective subset of MLP layers within the Transformer is expanded into MoE layers, with each expert initialized using the same MLP weights from the dense MLP layer of the dense CLIP model, while the router weights are randomly initialized. All other layers, such as attention and embedding layers, are copied directly from the original model. This approach ensures that only a sparse part of the model is modified, while most parameters remain unchanged, preserving the model’s core functionality. After initialization, the upcycled model undergoes additional training steps with the original hyperparameters like batch size, while slightly reducing the learning rate and weight decay to improve performance and stabilize training. The upcycle training process is illustrated in the right panel of figure 2.
Figure 3: Impact of LIMOE auxiliary loss on different training configurations. We observe adding LIMOE loss could sometimes lead to instability issue especially when the backbones are not shared, while our upcycling recipe are more robust.
4 Experiments
Datasets.
For training, we used the same paired image-text dataset for the initial dense CLIP checkpoint training, CLIP-UP, as well as the baseline dense CLIP model. The training data includes WIT-3000M (Wu et al., 2024) and DFN-5B (Fang et al., 2023). For evaluation, we used the standard ImageNet (Deng et al., 2009; Shankar et al., 2020) image classification task and COCO (Lin et al., 2014) and Flickr30K (Plummer et al., 2017) image-to-text and text-to-image retrieval tasks. The input image resolution is 224 for all of the datasets.
Setup.
We begin by training a dense CLIP model for 440k steps, which serves as the initial stage. Following this, we refine the dense model into an MoE version using the proposed upcycle training, then continuously training it for an additional 350k steps. Both stages use the AdamW optimizer with a batch size of 32k. For the dense model, we apply a learning rate of $5\times 10^{-4}$ and a weight decay of 0.2. During the upcycle training phase, the learning rate is reduced to $5\times 10^{-5}$, and the weight decay to 0.05. In the MoE model, half of the MLP layers in the Transformer are replaced with MoE layers, creating an alternating [dense, sparse] structure for both modalities. Each MoE layer includes 8 experts, with the router selecting the top 2 experts for activation. During training, we choose $\alpha=0.01$ and $\beta=0.001$ in the router auxiliary loss to ensure balanced expert distribution without dominating the primary cross-entropy objective. Hyper-parameters were fine-tuned based on preliminary experiments to improve performance and ensure training stability.
Table 2: Comparing performance of CLIP-UP and baseline models including LIMOE (Mustafa et al., 2022) across different model sizes. The sparse upcycling leverages the CLIP dense checkpoint trained for 440k steps. CLIP-UP is further trained for additional 350k steps (790k steps in total). We also train CLIP with 790k steps to have comparable training steps to CLIP-UP.
@@ 4 Experiments
Additionally, we train another dense CLIP model for a total of 790k steps to enable a fair comparison. This model is trained with the same configuration as mentioned earlier, using a learning rate of $5\times 10^{-4}$ and a weight decay of 0.2 to optimize performance.

### v2

@@ Abstract
CLIP-UP: A Simple and Efficient Mixture-of-Experts CLIP Training Recipe with Sparse Upcycling
Xinze Wang   Chen Chen   Yinfei Yang   Hong-You Chen   Bowen Zhang Aditya Pal   Xiangxin Zhu   Xianzhi Du Apple xinze_wang@apple.com
@@ 1 Introduction
CLIP Radford et al. (2021); Jia et al. (2021) has become foundational across domains such as image classification, multimodal retrieval, and AI-driven multimodality content generation Zhou et al. (2022); Rao et al. (2022); Gan et al. (2022); Ramesh et al. (2021); Liu et al. (2023). As applications grow, scaling CLIP becomes essential. Most efforts focus on enlarging dense models Cherti et al. (2023), which improves performance but incurs high computational and inference costs.
Figure 1: Our proposed MoE CLIP pre-training recipe. We highlight key factors for efficient training, including backbone sharing, training from scratch vs. sparse upcycling, and auxiliary losses. A detailed analysis is provided in Section 3.1 and Section 3.2.
An efficient alternative is sparse modeling with Mixture-of-Experts (MoE) Mustafa et al. (2022); Shazeer et al. (2017). However, training MoE-based CLIP models like LIMOE Mustafa et al. (2022) from scratch remains expensive and often requires auxiliary losses for stability. For instance, LIMOE outperforms dense CLIP but demands 1.35× more training FLOPs Mustafa et al. (2022).
To address this, we explore sparse upcycling Komatsuzaki et al. (2023), which initializes MoE layers from a pre-trained dense model. As shown in Figure 1, our extensive experiments demonstrate that sparse upcycling with a separated backbone achieves the best performance while reducing training ZFLOPs from 4.2 to 3.7 compared to training from scratch. Although LIMOE’s entropy losses Mustafa et al. (2022) improve shared-backbone models trained from scratch, they still underperform other setups. Section 3.1 details these strategies and the effects of auxiliary losses.
In contrast, we propose CLIP-UP, a single-stage sparse upcycling method for CLIP. By leveraging pre-trained weights, CLIP-UP provides a warm start that boosts efficiency and surpasses both dense continued training and sparse-from-scratch methods across model scales. [^1] Concurrent work CLIP-MoE Zhang et al. (2024) also explores MoE upcycling, using cluster-and-contrast learning to initialize experts. However, it requires additional training stages per expert, making it difficult to scale.
Our main contributions are:
@@
    We introduce CLIP-UP, a simple and effective training recipe for MoE CLIP models via sparse upcycling, avoiding complex auxiliary losses and outperforming existing methods across shared and separated backbones.
@@
    CLIP-UP significantly improves performance on text-image retrieval, surpassing dense CLIP by 7.2% and 5.5% (recall@1) on COCO and Flickr30K, respectively, with a B/16 backbone.
@@
    We demonstrate CLIP-UP’s scalability from B/32 to L/14 and provide insights into key factors and challenges to inform future design.
2 CLIP-UP
Figure 2 illustrates the CLIP-UP architecture and training strategy. We detail both in this section.
Figure 2: CLIP-UP overview with sparse upcycling initialization. Selected MLP layers are replaced with MoE layers, initialized from the dense checkpoint, while routers are randomly initialized.
@@ 2.1 CLIP
Given $n$ pairs of image and text captions $\{(\mathbf{I}_{j},\mathbf{T}_{j})\}_{j=1}^{n}$, CLIP Radford et al. (2021) learns image and text embeddings $f(\mathbf{I}_{j})$ and $g(\mathbf{T}_{j})$ using a contrastive loss. With batch size $B$ and temperature $\theta$, the loss is
$\mathcal{L}_{\text{Contrastive}}$ $=-\frac{1}{2B}\sum_{j=1}^{B}(\log\frac{e^{\text{sim}(f(\mathbf{I}_{j}),g(\mathbf{T}_{j}))/\theta}}{\sum_{k=1}^{B}e^{\text{sim}(f(\mathbf{I}_{j}),g(\mathbf{T}_{k}))/\theta}}$ (1)
$+\log\frac{e^{\text{sim}(f(\mathbf{I}_{j}),g(\mathbf{T}_{j}))/\theta}}{\sum_{k=1}^{B}e^{\text{sim}(f(\mathbf{I}_{k}),g(\mathbf{T}_{j}))/\theta}})$
@@ 2.2 CLIP with Mixture-of-Experts upcycling
Each MoE layer consists of $E$ MLP experts and a router that activates the top-$K$ experts per input token based on predicted gating logits. Let $\mathbf{X}_{j}\in\mathbb{R}^{D}$ be the input for the $j$-th token, $\mathbf{G}_{e,j}\in\mathbb{R}^{D}$ the gating logits, and $\mathbf{W}_{e}\in\mathbb{R}^{D}$ the router weights for expert $e$. The output $\text{MoE}(\mathbf{X}_{j})$ is computed as:
$\text{MoE}(\mathbf{X}_{j})=\mathbf{X}_{j}+\sum_{e\in\text{Top-K}}G_{e,j}\text{MLP}_{e}(\mathbf{X}_{j})$
$\hskip 14.22636ptG_{e,j}=\begin{cases}\text{Softmax}(\mathbf{W}_{e}^{T}\mathbf{X}_{j}),&\text{if}\ e\in\text{Top-K},\\ 0,&\text{otherwise},\end{cases}$
Each expert is assigned a fixed buffer capacity Fedus et al. (2022), allowing it to process a limited number of tokens at a time. With capacity factor $C$, batch tokens $B_{t}$, the capacity per expert is $B_{e}=(B_{t}/E)\times C$. This ensures computational efficiency and effective resource management. Tokens are assigned to experts on a "first-come-first-serve" basis Fedus et al. (2022). This simple mechanism avoids prioritization overhead while efficiently distributing tokens across experts.
Auxiliary loss. Simplified token assignment reduces overhead but risks imbalanced token distribution, leading to token dropping and performance degradation Zeng et al. (2024). To mitigate this, we adopt an auxiliary loss Zoph et al. (2022) combining load balance loss and router $z$-loss with scaling factors $\alpha$ and $\beta$. The load balance loss promotes uniform token allocation across experts. For a sequence of length $S$, it is defined as:
$\mathcal{L}_{\text{Balance}}=\alpha\cdot\sum_{e=1}^{E}R_{e}\cdot P_{e}$
@@
where $R_{e}=\frac{E}{K\cdot S}\sum_{j=1}^{S}\mathbbm{1}(\text{Token $j$}\rightarrow\text{Expert $e$})$ and $P_{e}=\frac{1}{S}\sum_{j=1}^{S}\mathbf{G}_{e,j}$ , denoting the token assignment ratio and average router probability for expert $e$ respectively.
@@
The router $z$-loss stabilizes gating by regularizing router logits to keep outputs within a reasonable range. It is defined as:
@@
$\mathcal{L}_{\text{Router}}=\beta\cdot\frac{1}{S}\sum_{j=1}^{S}\left(\log\sum_{e=1}^{E}e^{\mathbf{G}_{e,j}}\right)^{2}$
LIMOE auxiliary loss. We experimented with LIMOE’s local and global entropy losses Mustafa et al. (2022), tuning hyperparameters accordingly. While LIMOE auxiliary loss improves shared backbone trained from scratch, it underperforms in other settings. Therefore, we use load balance and router-$z$ losses as our auxiliary loss.
2.3 Sparse Upcycling Training
Sparse upcycling begins with a pre-trained dense CLIP, replacing selected MLP layers with MoE layers—experts initialized from the dense weights and routers randomly initialized. All other layers remain unchanged. The model is then fine-tuned with slightly reduced learning rate and weight decay for improved stability, as shown in Figure 2.
3 Experiments
Datasets.
We trained both the initial dense CLIP checkpoint, CLIP-UP, and the baseline model on the same paired image-text datasets—WIT-300M (Wu et al., 2024) and DFN-5B (Fang et al., 2023). Evaluation was performed on ImageNet (Deng et al., 2009; Shankar et al., 2020) for classification and on COCO (Lin et al., 2014) and Flickr30K (Plummer et al., 2017) for image-text retrieval, with additional benchmarks provided in the Appendix C.2. The input image resolution is 224 for all of the datasets.
@@ 3 Experiments
Table 1: Ablation study comparing shared vs. separated backbones and training from scratch vs. sparse upcycling, evaluated on ImageNet (Accuracy@1 %) and COCO/Flickr30K text-to-image (T2I)/image-to-text (I2T) retrieval (Recall@1 %)
@@
Table 2: Performance comparison of CLIP-UP, dense models, and LIMOE across model sizes. CLIP-UP is upcycled from a 440k-step CLIP checkpoint with 350k additional steps; a 790k-step dense CLIP is trained for fair comparison.
@@
Setup.
We train a dense CLIP model for 440k steps, then upcycle it into an MoE version with 350k additional steps. Both use AdamW with a 32k batch size; the dense model uses a learning rate of $5\times 10^{-4}$ and weight decay of 0.2, reduced to $5\times 10^{-5}$ and 0.05 for upcycling. In the MoE model, half of the Transformer’s MLP layers follow an alternating [dense, sparse] pattern Zoph et al. (2022); Du et al. (2022), each sparse layer using 8 experts with top-2 routing. Router loss coefficients $\alpha=0.01$ and $\beta=0.001$ balance expert usage without dominating training Zoph et al. (2022); Xue et al. (2024). For fair comparison, we also train a dense CLIP for 790k steps using the same settings.